# Remove Symmetries to Control Model Expressivity

### v1

@@ 6 Experiment
First, we show that the proposed method is compatible with standard training methods. We then apply the method to a few settings where symmetry is known to be a major problem in training. We see that the algorithm leads to improved model performance on these problems.
@@ 6.1 Compatibility with Standard Training
Let us first consider the classical problem of linear regression with $d$-dimensional data, where one wants to find $\min_{w}\sum_{i}(w^{T}x_{i}-y_{i})^{2}$. Here, the use of weight decay has a well-known effect of preventing the divergence of generalization loss at a critical dataset size $N=d$ [25, 26]. This is due to the fact that the Hessian matrix of the loss becomes singular exactly at $N=d$ (at infinite $N$ and $d$). The use of weight decay shifts all the eigenvalues of the Hessian by $\gamma$ and removes this singularity. In this case, one can show that the proposed method is essentially identical to the simple ridge regression. The ridge solution is $w^{*}=\mathbb{E}[\gamma I+A]^{-1}\mathbb{E}[xy]$, where $A=\mathbb{E}[xx^{T}]$, and the solution to the biased model is
@@
Figure 3: Training with syre in standard settings. The result shows that biasing the models by a small static bias does not change the performance of standard training settings. Left: Application of the method to a linear regression problem. Here, $\alpha^{-1}=d/N$ is the degree of parameterization. A well-known use of weight decay is to prevent double descent when $\alpha=1$. Here, we see that the proposed method works as well as vanilla weight decay. Because there is no reflection symmetry in the problem, the proposed method should approximate vanilla weight decay. Mid: Test accuracy of Resnet18 on the CIFAR-10 datasets. The blue line denotes the performance of Resnet, and the shadowed area denotes its standard deviation estimated over $10$ trials. For $\sigma_{0}<0.2$, there is no significant difference between the performance of the vanilla Resnet and syre Resnet. Right: linear regression with a redundant parametrization [27, 28]. The loss function takes the form $\ell(u,w)=\|(u\odot w)^{T}x-y\|^{2}$. Due to symmetry, the point $(u_{i},w_{i})=0$ is a low-capacity state where the $i$-th neuron is “dead”. Training with style, the model stayed away from any trapping low-capacity state during training. In comparison, training with vanilla SGD or a heuristic for fixing the weights does not fix the problem of collapsing to a low-capacity state.
@@
A minimal model with emergent interest in the problem of compressing neural networks is the reparametrized version of linear regression [28], the loss function of which is $\ell(u,w)=\|(u\odot w)^{T}x-y\|^{2}$, where we let $u,\ w,\ x\in\mathbb{R}^{200}$ and $y\in\mathbb{R}$. Due to the rescaling symmetry between every parameter $u_{i}$ and $w_{i}$, the solutions where $u_{i}=w_{i}=0$ is a low-capacity state where the $i$-th neuron is “dead.” For this problem, we compare the training with standard SGD and syre. We also compare with a heuristic method (W-fix), where a fraction $\phi=0.3$ of weights of every layer is held fixed with a fixed variance $\kappa=0.01$ at initialization. This method has been suggested in [16] as a heuristic for removing symmetries and is found to work well when there is permutation symmetry. We see that both the vanilla training and the W-fix collapse to low-capacity states during training, whereas the proposed method stayed away from them throughout. The reason is that the proposed method is model-independent and symmetry-agnostic, working effectively for any type of possibly unknown symmetry in an arbitrary architecture.
@@
We also benchmark the performance of the proposed method for ResNet18 with different $\sigma_{0}$ on the CIFAR-10 datasets in Figure 3. When $\sigma_{0}=0$, the syre model is equivalent to the vanilla model. Figure 3 shows that the difference in performance between the vanilla Resnet and syre Resnet is very small and becomes neglectable when $\sigma_{0}<0.2$. We thus recommend a default value of $\sigma_{0}\leq 0.01/\sqrt{d}$.
@@ 6.2 Benchmarking Symmetry Removal
In this section, we benchmark the effect of symmetry control of the proposed method for two controlled experiments. To compare the influence of syre and other training methods on the degree of symmetry, we consider minimizing the following objective function
$(w^{T}w)^{2}-w^{T}Bw:=(w^{T}w)^{2}-\sum_{i=1}^{d}\lambda_{i}(v_{i}^{T}w)^{2},$
Benchmarking where $w\in\mathbb{R}^{d}$ is the optimization parameter and $B\in\mathbb{R}^{d\times d}$ is a given symmetric matrix with eigenvalues $\lambda_{i}$ and eigenvectors $v_{i}$ ($v_{i}^{T}v_{i}=1$). The objective function has $n$ reflection symmetries $P_{i}w:=w-2(v_{i}^{T}w)v_{i}$. Hence, we define the degree of symmetry as $\sum_{i=1}^{d}\mathbf{1}\{v_{i}^{T}w<c_{\rm th}\}$, where $c_{\rm th}$ is a given threshold. Depending on the spectrum of $B$, the nature of the task is different. We thus consider two types of spectra: (1) an unstructured spectrum where $B=G+G^{T}$ for a Gaussian matrix $G$, and (2) a structured spectrum where $B={\rm diag}(v)$ where $v$ is a random Gaussian vector. Conceptually, the first type is more similar to rotation and double rotation symmetries in neural networks where the basis can be arbitrary, while the second is a good model for common discrete symmetries where the basis is often diagonal or sparse. For the first case we choose $c_{\rm th}=10^{-3}$ and for the second case we choose $c_{\rm th}=10^{-1}$.
@@ 6.3 Feature and Neuron Collapses in Supervised Learning
Figure 5: Performance of a 4-layer FCN for datasets with various weight decay $\gamma$ and data distributions with varyings strengths of linear correlation $\alpha$. As the theory predicts, the covariance of the vanilla model has a low-rank structure and performs significantly worse. In the main figures, solid lines denote training accuracy and dashed lines denote test accuracy. The dashed black line corresponds to random guess. Subfigures show the rank of the covariance matrix of the first layer output before (solid lines) and after (dashed lines) activation with batch size $1000$ scaled by $\alpha^{2}$. We set eigenvalues smaller than $10^{-4}$ to 00. Left: $\alpha=1$ and different $\gamma$. Right: $\gamma=0.01$ and different $\alpha$.
See Figure 5, where we train the vanilla and syre four-layer networks with various levels of weight decay $\gamma$ and various levels of input-output covariance $\alpha$. The dataset is constructed by rescaling the input by a factor of $\alpha$ for the MNIST dataset. The theory predicts that the syre model can remove the permutation symmetry in the hidden layer. This is supported by subfigures in Figure 5, where vanilla training results in a low-rank solution. Meanwhile, the accuracy of the low-rank solution is significantly lower for a large $\gamma$ or a small $\alpha$, which corresponds to the so-called neural collapses. Also, we observe that syre shifts the eigenvalues of the representation by a magnitude proportional to $\sigma_{0}$, thus explaining the robustness of the method against collapses in the latent representation (See Figure 11).
@@ 6.4 Posterior Collapse in Bayesian learning
Ref. [29] points out that a type of posterior collapse in Bayesian learning [30, 31] is caused by the low-rankness of the solutions. We show that training with syre could overcome this kind of posterior collapse. In Figure 6, we train a $\beta$-VAE [32, 33] on the Fashion MNIST dataset. Following Ref. [29], we use $\beta$ to weigh the KL loss, which can be regarded as the strength of prior matching. Both the encoder and the decoder are a two-layer network with SiLU activation. The hidden dimension and the latent dimension are $200$. Only the encoder has weight decay because the low-rank problem is caused by the encoder rather than the decoder. We also choose the prior variance of the latent variable to be $\eta_{enc}=0.01$. Other settings are the same as [29]. Posterior collapse happens at $\beta=10$, signalized by a large reconstruction loss in the right side of Figure 6. However, the reconstruction loss decreases, and the rank of the encoder output increases (according to the left side of Figure 6) after we use weight decay and syre. This is further verified by the generated image in Figure 7. Therefore, we successfully remove the permutation symmetry of the encoder.
Figure 6: Rank and reconstruction loss for a VAE is trained on the Fashion MNIST dataset. The covariance of the vanilla model has a low-rank structure and larger reconstruction loss. More importantly, posterior collapse happens at $\beta=5$ but is mitigated with weight decay. Left: the rank of the encoder output with batch size $1000$. We set eigenvalues smaller than $10^{-6}$ to 00. Right: reconstruction loss of vanilla and syre models.
@@ 6.5 Low-Capacity Trap in Self-supervised Learning
A common but bizarre practice in self-supervised learning (SSL) is to throw away the last layer of the trained model and use the penultimate learning representation, which is found to have much better expressiveness than the last layer representation. From the perspective of symmetry, this problem is caused by the rotation symmetry of the last weight matrix in the SimCLR loss [8]. We train a Resnet-18 together with a two-layer projection head over the CIFAR-100 dataset according to the setting for training SimCLR in [34]. Then, a linear classifier is trained using the learned representations. Our implementation reproduces the typical accuracy of SimCLR over the CIFAR-100 dataset [35]. As in [34], the hidden layer before the projection head is found to be a better representation than the layer after. Therefore, we apply our syre method to the projection head or to all layers. According to Table 1, syre removes the low-rankness of the learned features and increases the accuracy trained with the features after projection while not changing the representation before projection. Thus, symmetry-induced reduction in model capacity can explain about $50\%$ of the performance difference between the representation of the two layers. Also, an interesting observation is that just improving the expressivity of the last layer is insufficient to close the gap between the performance of the last layer and the penultimate layer. This helps us gain a new insight: symmetry is not the only reason why the last layer representation is defective.
Figure 8: Performance and accuracy of a CNN trained on a continual learning task (permuted MNIST [36, 37]). The main figure shows the test accuracy, and the subfigure shows the rank of the convolution layers output with batch size $1000$, where we set eigenvalues smaller than $10^{-4}$ to 00. The results suggest that the rank of the vanilla model gradually decreases, and the model completely collapses after the sixth task, while the syre model remains unaffected.
Figure 9: Loss of plasticity in continual learning in an RL setting. We use the PPO algorithm [38] to solve the Slippery-Ant problem [39]. The rank and the performance of the vanilla PPO decrease quickly, while the rank and the performance of PPO with syre remain the same, beyond that of PPO with weight decay. Left: the effective rank of the policy network as defined in [39]. Right: returns. Each trajectory is averaged over $5$ different random seeds.
@@ 6.6 Loss of Plasticity in Continual Learning
A form of low-capacity collapse also happens during continual learning, i.e., the plasticity of the network gradually decreases as the model is trained on more and more tasks. This problem is common in both supervised and reinforcement learning settings and may also be relevant to the finetuning of large language models [40, 41, 42, 43].
@@
In Figure 9, we use the PPO algorithm for the Slippery-Ant problem [39], a continual variant of the Pybullet’s Ant problem [44] with friction that changes every $5M$ steps. Hyperparameters for the PPO algorithm are borrowed from [39], and we use a weight decay of $0.002$ for both PPO with weight decay and with syre. Figure 9 suggests that syre is effective in maintaining the rank of the model during continual training and obtains better performance than pure weight decay.
@@ 7 Conclusion
Symmetry-induced neural-network training problems exist extensively in machine learning. We have shown that the existence of symmetries in the model or loss function may severely limit the expressivity of the trained model. We then developed a theory that leverages the power of representation theory to show that adding random static biases to the model, along with weight decay, is sufficient to remove almost all symmetries, explicit or hidden. We have demonstrated the relevance of the method to a broad range of applications in deep learning, and a possible future direction is to deploy the method in large language models, which naturally contain many symmetries. Lastly, it is worth noting that on its own, symmetry is neither good nor bad. For example, practitioners may be interested in introducing symmetries to the model architecture in order to control the capacity of the model [28]. However, with too much symmetry, the training of models becomes slow and likely to contain many low-capacity traps. Meanwhile, a model completely without symmetry may have undesirably high capacity and be more prone to overfitting. Having the right degree of symmetry might thus be crucial for achieving both smooth optimization and good generalization. With our proposed method, it becomes increasingly possible to deliberately fine-grain engineer symmetries in the loss function, introducing desired symmetries and removing undesirable ones.
Acknowledgements
The authors acknowledge support from NTT Research, and from the Institute for Artificial Intelligence and Fundamental Interactions (IAIFI) through NSF Grant No. PHY-2019786.
@@ A.3 Proof of Proposition 3
According to Ref. [11, Theorem 4], we have
@@ B.2 Supervised Learning
Figure 11: The spectrum of the covariance matrix of the vanilla model (Left) and the syre model (Right) for $\gamma=0.01$ and $\alpha=1$. Clearly, the vanilla model learns a low-rank solution.

### v3

@@ 6 Experiment
First, we show that the proposed method is compatible with standard training methods. We then apply the method to a few settings where symmetry is known to be a major problem in training. We see that removing symmetries with the proposed method is well correlated with improved model performance for these problems. Additional experiments are presented in Appendix B, including the time and memory usage of syre and its application to transformers and improving mode connectivity. [^6] An implementation of syre can be found at https://github.com/xu-yz19/syre/.
@@ 6.1 Compatibility with Standard Training
Let us first consider the classical problem of linear regression with $d$-dimensional data, where one wants to find $\min_{w}\sum_{i}(w^{T}x_{i}-y_{i})^{2}$. Here, the use of weight decay has a well-known effect of preventing the divergence of generalization loss at a critical dataset size $N=d$ (Krogh & Hertz, 1992; Hastie et al., 2019). This is due to the fact that the Hessian matrix of the loss becomes singular exactly at $N=d$ (at infinite $N$ and $d$). The use of weight decay shifts all the eigenvalues of the Hessian by $\gamma$ and removes this singularity. In this case, one can show that the proposed method is essentially identical to the simple ridge regression. The ridge solution is $w^{*}=\mathbb{E}[\gamma I+A]^{-1}\mathbb{E}[xy]$, where $A=\mathbb{E}[xx^{T}]$, and the solution to the biased model is
@@
Figure 3: Training with syre in standard settings. The result shows that biasing the models by a small static bias does not change the performance of standard training settings. Left: Application of the method to a linear regression problem. Here, $\alpha^{-1}=d/N$ is the degree of parameterization. A well-known use of weight decay is to prevent double descent when $\alpha=1$. Here, we see that the proposed method works as well as vanilla weight decay. Because there is no reflection symmetry in the problem, the proposed method should approximate vanilla weight decay. Mid: Test accuracy of Resnet18 on the CIFAR-10 datasets. The blue line denotes the performance of Resnet, and the shadowed area denotes its standard deviation estimated over $10$ trials. For $\sigma_{0}<0.2$, there is no significant difference between the performance of the vanilla Resnet and syre Resnet. Right: linear regression with a redundant parametrization (Poon & Peyré, 2021). The loss function takes the form $\ell(u,w)=\|(u\odot w)^{T}x-y\|^{2}$. Due to symmetry, the point $(u_{i},w_{i})=0$ is a low-capacity state where the $i$-th neuron is “dead”. Training with style, the model stayed away from any trapping low-capacity state during training. In comparison, training with vanilla SGD or a heuristic for fixing the weights does not fix the problem of collapsing to a low-capacity state.
@@
A minimal model with emergent interest in the problem of compressing neural networks is the reparametrized version of linear regression (Poon & Peyré, 2021; Ziyin & Wang, 2023), the loss function of which is $\ell(u,w)=\|(u\odot w)^{T}x-y\|^{2}$, where we let $u,\ w,\ x\in\mathbb{R}^{200}$ and $y\in\mathbb{R}$. Due to the rescaling symmetry between every parameter $u_{i}$ and $w_{i}$, the solutions where $u_{i}=w_{i}=0$ is a low-capacity state where the $i$-th neuron is “dead.” For this problem, we compare the training with standard SGD and syre. We also compare with a heuristic method (W-fix), where a fraction $\phi=0.3$ of weights of every layer is held fixed with a fixed variance $\kappa=0.01$ at initialization. This method has been suggested in Lim et al. (2024) as a heuristic for removing symmetries and is found to work well when there is permutation symmetry. We see that both the vanilla training and the W-fix collapse to low-capacity states during training, whereas the proposed method stayed away from them throughout. The reason is that the proposed method is model-independent and symmetry-agnostic, working effectively for any type of possibly unknown symmetry in an arbitrary architecture.
@@
We also benchmark the performance of the proposed method for ResNet18 with different $\sigma_{0}$ on the CIFAR-10 datasets in Figure 3. When $\sigma_{0}=0$, the syre model is equivalent to the vanilla model. Figure 3 shows that the difference in performance between the vanilla Resnet and syre Resnet is very small and becomes neglectable when $\sigma_{0}<0.2$.
@@ 6.2 Benchmarking Symmetry Removal
In this section, we benchmark the effect of symmetry control of the proposed method for two controlled experiments. To compare the influence of syre and other training methods on the degree of symmetry, we consider minimizing the following objective function:
$\vspace{-2mm}(w^{T}w)^{2}-w^{T}Bw:=(w^{T}w)^{2}-\sum_{i=1}^{d}\lambda_{i}(v_{i}^{T}w)^{2},$
where $w\in\mathbb{R}^{d}$ is the optimization parameter and $B\in\mathbb{R}^{d\times d}$ is a given symmetric matrix with eigenvalues $\lambda_{i}$ and eigenvectors $v_{i}$ ($v_{i}^{T}v_{i}=1$). The objective function has $n$ reflection symmetries $P_{i}w:=w-2(v_{i}^{T}w)v_{i}$. Hence, we define the degree of symmetry as $\sum_{i=1}^{d}\mathbf{1}\{v_{i}^{T}w<c_{\rm th}\}$, where $c_{\rm th}$ is a given threshold. Depending on the spectrum of $B$, the nature of the task is different. We thus consider two types of spectra: (1) an unstructured spectrum where $B=G+G^{T}$ for a Gaussian matrix $G$, and (2) a structured spectrum where $B={\rm diag}(v)$ where $v$ is a random Gaussian vector. Conceptually, the first type is more similar to rotation and double rotation symmetries in neural networks where the basis can be arbitrary, while the second is a good model for common discrete symmetries where the basis is often diagonal or sparse. For the first case we choose $c_{\rm th}=10^{-3}$ and for the second case we choose $c_{\rm th}=10^{-1}$.
@@ 6.3 Feature and Neuron Collapses in Supervised Learning
Figure 5: Performance of a 4-layer FCN for datasets with various weight decay $\gamma$ and data distributions with varyings strengths of linear correlation $\alpha$. As the theory predicts, the covariance of the model with vanilla weight decay has a low-rank structure and performs significantly worse. In the main figures, solid lines denote training accuracy and dashed lines denote test accuracy. The dashed black line corresponds to random guess. Subfigures show the rank of the covariance matrix of the first layer output before (solid lines) and after (dashed lines) activation with batch size $1000$. We set eigenvalues smaller than $10^{-4}$ to 00. Left: $\alpha=1$ and different $\gamma$. Right: $\gamma=0.01$ and different $\alpha$. A similar experiment for vision transformers is presented in Section B.5.
See Figure 5, where we train the vanilla and syre four-layer networks with various levels of weight decay $\gamma$ and input-output covariance $\alpha$. The dataset is constructed by rescaling the input by a factor of $\alpha$ for the MNIST dataset. The theory predicts that the syre model can remove the permutation symmetry in the hidden layer. This is supported by subfigures in Figure 5, where vanilla training results in a low-rank solution. Meanwhile, the accuracy of the low-rank solution is significantly lower for a large $\gamma$ or a small $\alpha$, which corresponds to the so-called neural collapses. Also, we observe that syre shifts the eigenvalues of the representation by a magnitude proportional to $\sigma_{0}$, thus explaining the robustness of the method against collapses in the latent representation (See Figure 8).
@@ 6.4 Posterior Collapse in Bayesian learning
Wang & Ziyin (2022) points out that a type of posterior collapse in Bayesian learning (Lucas et al., 2019; Wang et al., 2021) is caused by the low-rankness of the solutions. We show that training with syre could overcome this kind of posterior collapse. In Figure 9, we train a $\beta$-VAE (Kingma & Welling, 2013; Higgins et al., 2016) on the Fashion MNIST dataset. Following Wang & Ziyin (2022), we use $\beta$ to weigh the KL loss, which can be regarded as the strength of prior matching. Both the encoder and the decoder are a two-layer network with SiLU activation. The hidden dimension and the latent dimension are $200$. Only the encoder has weight decay because the low-rank problem is caused by the encoder rather than the decoder. We also choose the prior variance of the latent variable to be $\eta_{enc}=0.01$. Other settings are the same as Wang & Ziyin (2022). Posterior collapse happens at $\beta=10$, signalized by a large reconstruction loss in the right side of Figure 9. However, the reconstruction loss decreases, and the rank of the encoder output increases (according to the left side of Figure 9) after we use weight decay and syre. This is further verified by the generated image in Figure 10. Therefore, we successfully remove the permutation symmetry of the encoder.
@@ 6.5 Low-Capacity Trap in Self-supervised Learning
A common but bizarre practice in self-supervised learning (SSL) is to throw away the last layer of the trained model and use the penultimate learning representation, which is found to have much better expressiveness than the last layer representation. From the perspective of symmetry, this problem is caused by the rotation symmetry of the last weight matrix in the SimCLR loss. We train a Resnet-18 together with a two-layer projection head over the CIFAR-100 dataset according to the setting for training SimCLR in Chen et al. (2020). Then, a linear classifier is trained using the learned representations. Our implementation reproduces the typical accuracy of SimCLR over the CIFAR-100 dataset (Patacchiola & Storkey, 2020). As in Chen et al. (2020), the hidden layer before the projection head is found to be a better representation than the layer after. Therefore, we apply our syre method to the projection head or to all layers. According to Table 1, syre removes the low-rankness of the learned features and increases the accuracy trained with the features after projection while not changing the representation before projection. Thus, symmetry-induced reduction in model capacity can explain about $50\%$ of the performance difference between the representation of the two layers. Also, an interesting observation is that just improving the expressivity of the last layer is insufficient to close the gap between the performance of the last layer and the penultimate layer. This helps us gain a new insight: symmetry is not the only reason why the last layer representation is defective.
Figure 6: Loss of plasticity in continual learning in an RL setting. We use the PPO algorithm (Schulman et al., 2017) to solve the Slippery-Ant problem (Dohare et al., 2023). The rank and the performance of the vanilla PPO decrease quickly, while the rank and the performance of PPO with syre remain the same, beyond that of PPO with weight decay. Left: the effective rank of the policy network as defined in Dohare et al. (2023). Right: returns. Each trajectory is averaged over $5$ different random seeds.
@@ 6.6 Loss of Plasticity in Continual Learning
A form of low-capacity collapse also happens during continual learning, i.e., the plasticity of the network gradually decreases as the model is trained on more and more tasks. This problem is common in both supervised and reinforcement learning settings and may also be relevant to the finetuning of large language models (McCloskey & Cohen, 1989; Ash & Adams, 2020; Dohare et al., 2021; Abbas et al., 2023).
Supervised Learning.
@@
In Figure 6, we use the PPO algorithm for the Slippery-Ant problem (Dohare et al., 2023), a continual variant of the Pybullet’s Ant problem (Coumans & Bai, 2016) with friction that changes every $5M$ steps. Hyperparameters for the PPO algorithm are borrowed from Dohare et al. (2023), and we use a weight decay of $0.002$ for both PPO with weight decay and with syre. Figure 6 suggests that syre is effective in maintaining the rank of the model during continual training and obtains better performance than pure weight decay.
@@ 7 Conclusion
Symmetry-induced neural-network training problems exist extensively in machine learning. We have shown that the existence of symmetries in the model or loss function may severely limit the expressivity of the trained model. We then developed a theory that leverages the power of representation theory to show that adding random static biases to the model, along with weight decay, is sufficient to remove almost all symmetries, explicit or hidden. We have demonstrated the relevance of the method to a broad range of applications in deep learning, and a possible future direction is to deploy the method in large language models, which naturally contain many symmetries. Lastly, it is worth noting that on its own, symmetry is neither good nor bad. For example, practitioners may be interested in introducing symmetries to the model architecture in order to control the capacity or hierarchy of the model (Ziyin et al., 2025). However, with too much symmetry, the training of models becomes slow and likely to contain many low-capacity traps. Meanwhile, a model completely without symmetry may have undesirably high capacity and be more prone to overfitting. Having the right degree of symmetry might thus be crucial for achieving both smooth optimization and good generalization. With our proposed method, it becomes increasingly possible to deliberately fine-grain engineer symmetries in the loss function, introducing desired symmetries and removing undesirable ones.
@@ A.3 Proof of Proposition 3
According to  Ziyin (2024, Theorem 4), we have
@@ B.2 Supervised Learning
Figure 8: The spectrum of the covariance matrix of the model with vanilla weight decay (Left) and the syre model (Right) for $\gamma=0.01$ and $\alpha=1$. Clearly, the vanilla model learns a low-rank solution.
B.3 Posterior Collapse
See Figures 9 and 10.
Figure 9: Rank and reconstruction loss for a VAE is trained on the Fashion MNIST dataset. The covariance of the model with vanilla weight decay has a low-rank structure and larger reconstruction loss. More importantly, posterior collapse happens at $\beta=5$ but is mitigated with weight decay. Left: the rank of the encoder output with batch size $1000$. We set eigenvalues smaller than $10^{-6}$ to 00. Right: reconstruction loss of vanilla and syre models.
B.4 Continual Learning
See Figure 11.
Figure 11: Performance and accuracy of a CNN trained on a continual learning task (permuted MNIST (Goodfellow et al., 2013; Kirkpatrick et al., 2017)). The main figure shows the test accuracy, and the subfigure shows the rank of the convolution layers output with batch size $1000$, where we set eigenvalues smaller than $10^{-4}$ to 00. The results suggest that the rank of the model with vanilla weight decay gradually decreases, and the model completely collapses after the sixth task, while the syre model remains unaffected.
Figure 12: Test accuracy and rank of ViT for CIFAR10 (adapted from Yoshioka (2024)) with various weight decay $\gamma$ and data distributions with varyings strengths of linear correlation $\alpha$. As the theory predicts, the covariance of features of the model with vanilla weight decay has a low-rank structure and performs significantly worse. Dashed black lines correspond to random guesses. Subfigures show the rank of the covariance matrix of the final layer input with batch size $512$. We set eigenvalues smaller than $10^{-4}$ to 00. Left: $\alpha=1$ and different $\gamma$. Right: $\gamma=0.0001$ and different $\alpha$.
B.5 Vision Transformer (ViT)
Similar to Section 6.3, we train the vanilla and syre ViT with various levels of weight decay $\gamma$ and various levels of input-output covariance $\alpha$. The dataset is constructed by rescaling the input by a factor of $\alpha$ for the CIFAR10 dataset. As the theoretical prediction, we can see that syre removes the rotation symmetries symmetry of transformers (Kobayashi et al., 2024) and leads to a full-rank solution with higher accuracy, as shown in Figure 12. Moreover, syre also performs better than the W-asymmetric and $\sigma$-asymmetric networks proposed in Lim et al. (2024). We implement $\sigma$-asymmetric networks using the recommended hyperparameters in Lim et al. (2024), and implement W-asymmetric networks by setting the same hyperparameters for each layer and doing a grid search.
B.6 Time and Memory Consumption
The time and memory consumption of various models used in previous sections is given in Tables 2 and 3. In all experiments, we train the models on the CIFAR10 dataset with a signal A6000 GPU. It is clear that syre has a neglectable influence on time and memory consumption. This is as expected because syre only adds a static bias to training.
batchsize MLP syre-MLP ResNet syre-ResNet ViT syre-ViT
512 $95\pm 0.1$ $97\pm 1$ $18.4\pm 0.7$ $19.2\pm 0.9$ $80\pm 1$ $84\pm 2$
256 $48\pm 1$ $49\pm 1$ $18.6\pm 0.2$ $19.6\pm 0.4$ $48\pm 5$ $57\pm 3$
128 $26\pm 1$ $26\pm 1$ $19.9\pm 0.4$ $20.9\pm 0.3$ $39\pm 1$ $47\pm 2$
Table 2: The per-batch time (ms) of various models.
batchsize MLP syre-MLP ResNet syre-ResNet ViT syre-ViT
512 382 384 7020 7049 4216 4218
256 360 382 3936 4254 2482 2484
128 360 382 2096 2190 1574 1574
Table 3: The GPU memory usage (M) of various models.
B.7 Linear Model Connectivity
Figure 13: Linear mode connectivity: test accuracy and loss curves along linear interpolations between trained networks. We train MLPs over the MNIST dataset.
We also test the influence of syre on the linear model connectivity with the same setting as Lim et al. (2024). Specifically, we obtain two well-behaved MLP with parameters $\theta_{1}$ and $\theta_{2}$, and test the performance of another MLP with parameters $\lambda\theta_{1}+(1-\lambda)\theta_{2}$ for $10<\lambda<$ 1. By removing the permutation symmetry of MLP, we expect that the MLP with parameters $\lambda\theta_{1}+(1-\lambda)\theta_{2}$ also performs well. Figure 13 suggests that syre performs similarly to W-asymmetric networks proposed in Lim et al. (2024), while our methods have much fewer hyperparameters.
Appendix C Symmetry and Rank of the NTK
In this section, we show that the rank of the NTK and gradient reduces as more and more of the neurons are in the symmetric state (namely, the fixed point induced by the group averaging projector).
Due to the difficulty in computing the spectrum of the NTK, we restrict to a two-layer subnetwork within a large network trained on MNIST with the number of neurons $10\to 64\to 10$, with $1280$ many parameters. This means that the gradient second moment has dimension $1280\times 1280$, and the model is trained on $1500$ data points. Thus, the empirical NTK can be seen as a matrix having dimension $1500\times 1500$. Here, the symmetry projection we consider is due to the permutation symmetry, whose projection has a rank of $20$ (because every neuron has ten outgoing weights and incoming weights). Thus, according to the Propositions 3 and 2, the maximum possible rank for the NTK and the gradient is $1280-20\times n$, where $n$ is the number of neurons projected to the symmetry fixed point. See Figure 14.
Figure 14: Ranks of the gradient covariance and NTK (Left: beginning of training. Right: end of training.) against the theoretical upper bound implied by Propositions 3 and 2. Two observations agree with our expectation: (1) the ranks of both matrices decrease as more and more neurons are stuck at the entrapped subspace, and (2) the ranks are always upper bound by the theoretical upper bound.